solve for x and y in the equation:2^(x+4y)=1 and 2^(x+8y)=1/4
Answer by htmentor(789)

solve for x and y in the equation:2^(x+4y)=1 and 2^(x+8y)=1/4
==================
Take the log of both sides:
(x+4y)log 2 = log 1
(x+8y)log 2 = log(1/4)
The 1st equation gives x = -4y
Subst. into the 2nd:
(4y)log 2 = log(1/4)
y = -1/2
So x = -4*(-1/2) = 2
Ans: x=2, y=-1/2